5.1 Refer to Exercise 1.6, and derive the moment estimate of θ. Also, compare it with the MLE θ= 1/x. 1.6 If the independent r.v.s. Xi, .., X have the Geometric p.d.f.
Answer #1

p.d.f.of Geometric Distribution is given by:

f(z: θ) = 9(1-0)1-1,

for x = 1, 2,...

The mean of geometric distribution is given by:

E(X)=sum_{x=1}^{infty }xf(x)=sum_{x=1}^{infty }x heta (1- heta )^{x-1}

-e+29(1-0)+36(1-0)2 +

= heta (1+2 (1- heta )+3(1- heta )^{2}+..)

= heta (1- (1- heta ))^{-2},

since 1+2x + 32.2 + = (1-1)

= heta ( heta )^{-2}

=rac{1}{ heta }

Thus, we get:

E(X) = (1)

Given:

Independent r.v.'s: X1, X2,...,Xn.

So,

Mean is given by:

E(X)=ar{X}=rac{sum_{i=1}^{n}X_{i}}{n} (2)

Equating (1) & (2), we get:

ar{X}=rac{1}{ heta }

Thus, the moment estimate of heta is given by:

θ=.

This is same as MLE of

θ=